OpenStudy (anonymous):

A spring hanging from the ceiling vibrates up and down. it's vertical position at time (t) is given by f(t)=4 sin(3t). (a) Find the velocity of the spring at time t. (b) What is the spring's maximum speed. (c) what is location when it reaches its maximum speed.

OpenStudy (anonymous):

You know that the position as a function of time is: \[x(t) = 4\sin(3t)\] you also know that velocity is defined as: \[\frac{dx}{dt}\]

OpenStudy (anonymous):

can you find the derivative of: \[x(t) = 4\sin(3t)\]

OpenStudy (anonymous):

Yes that is the ans of a.what about b and c parts?

OpenStudy (anonymous):

so we want to find the maximum of: \[12\cos(3t)\]

OpenStudy (anonymous):

So this basically comes down to finding where the cosine function has it's maximums

OpenStudy (anonymous):

and then using the value for the maximum of a cosine in the equation

OpenStudy (anonymous):

So, what is the maximum of a cosine function?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yes, so the maximum value for the speed must be:

OpenStudy (anonymous):

i think 12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

At what time does the speed become 12?

OpenStudy (anonymous):

don't know.may be 0

OpenStudy (anonymous):

Yes, because cos(x) = 1 when x = 0

OpenStudy (anonymous):

Now that you know the time at which it reaches its maximum speed, you can plug that into the first equation

OpenStudy (anonymous):

which first equation?

OpenStudy (anonymous):

the equation for position as a function of time

OpenStudy (anonymous):

f(t)=4sin(3t)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what i do with this now? please explain

OpenStudy (anonymous):

i put t=0 so position is zero?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok.Thanks a lot.

OpenStudy (anonymous):

You are welcome